# Freezing-in the Axiverse

### v1

@@ 1 Introduction
The Strong Charge-Parity (CP) problem in the Standard Model (SM) is the statement that the strong interactions appear to conserve CP symmetry while the weak interactions nearly maximally violate CP. Indeed, the Strong CP-violating parameter $\bar{\theta}$ is constrained to be $|\bar{\theta}|\lesssim 10^{-10}$ Abel:2020pzs, while one would have expected $\bar{\theta}$ to be an ${\cal O}(1)$ number between $\pi$ to $\pi$, since a vanishing $\bar{\theta}$ does not lead to any enhanced symmetry of the SM. The existence of the quantum chromodynamics (QCD) axion is a hypothesized solution to this fine tuning problem. The QCD axion couples to gluons in the same way as how $\bar{\theta}$ appears in the SM Lagrangian. The presence of the QCD axion leads to a dynamical relaxation of (the effective) $\bar{\theta}$ to a vanishing value as the QCD axion gets localized to the minimum of its potential, developed during the QCD phase transition Peccei:1977hh; Peccei:1977ur; Weinberg:1977ma; Wilczek:1977pj. The QCD axion may also make up the dark matter (DM) of the Universe Preskill:1982cy; Abbott:1982af; Dine:1982ah.
@@
In minimal models, the QCD axion arises as the pseudo-Nambu-Goldstone boson of a global U(1) symmetry Kim:1979if; Shifman:1979if; Dine:1981rt; Zhitnitsky:1980tq. In such cases, the QCD axion solution to the Strong CP problem may not be robust. Planck scale-suppressed higher dimensional operators can break the global $U(1)$ symmetry and modify the QCD axion potential significantly, altering the locations of the minima of the axion potential. As a result, the QCD axion can get localized to a new minimum which does not correspond to a vanishing value of $\bar{\theta}$ and the Strong CP problem would reappear. This is known as the ‘axion quality problem’ Georgi:1981pu; Dine:1986bg; Kamionkowski:1992mf; Holman:1992us; Barr:1992qq; Ghigna:1992iv, since a robust solution to the Strong CP problem demands a ‘high quality’ symmetry that is robust to various symmetry breaking effects.
@@
Models of the extra-dimensional QCD axion typically have such a high quality symmetry. A classic example of this is the 5D $U(1)$ gauge field where the axion arises from the fifth component of the gauge field. The higher dimensional gauge invariance Choi:2003wr; Arkani-Hamed:2003xts, or an electric one-form symmetry of the extra-dimensional theory Reece:2025thc; Craig:2024dnl, provides a more robust protection against symmetry breaking effects. Contributions from symmetry breaking effects lead to exponentially-small corrections, in contrast to polynomial corrections found in minimal models. This mechanism also plays out in string theory, where a large number $\cal N$} (as many as 491, found in Type IIB string theory analysis in Demirtas:2018akl; Cicoli:2012sz) of pseudoscalars appear naturally as zero-modes of higher-dimensional gauge fields upon dimensional reduction. This leads to the notion of an ‘Axiverse’ Witten:1984dg; Choi:1985je; Barr:1985hk; Svrcek:2006yi; Arvanitaki:2009fg – a universe with many axions.
@@
The QCD axion is extremely light with mass $m_{a}\sim 6\,\mu{\rm eV}\,(f_{a}/10^{12}\,{\rm GeV})$ GrillidiCortona:2015jxo where $f_{a}$ is the axion decay constant that controls the axion coupling with the SM, with astrophysical constraints requiring $f_{a}\gtrsim 10^{9}$ GeV ParticleDataGroup:2024cfk. Interactions with the SM in the early universe may produce a “cosmic axion background” Turner:1986tb; Masso:2002np; Hannestad:2005df; Graf:2010tv; Salvio:2013iaa; Ferreira:2018vjj; Dror:2021nyr; DEramo:2021psx; DEramo:2021lgb; Dunsky:2022uoq; Langhoff:2022bij, which can still be around as relativistic radiation today. The axion energy density $\rho_{a}$, in comparison with the photon energy density $\rho_{\gamma}$, can be parametrized by the $\Delta N_{\rm eff}$ parameter,
@@
While the QCD axion alone contributes a small amount to $\Delta N_{\rm eff}$, one might expect the multitude of axions from the axiverse to give rise to a $\Delta N_{\rm eff}$ that is too large and ruled out by current the CMB and big-bang nucleosynthesis constraint, $\Delta N_{\rm eff}\lesssim 0.3$ Planck:2018vyg at 95% confidence. That is, assuming each axion attains a thermal abundance, we have $\Delta N_{\rm eff}\approx 0.027{\cal N}$. A scenario with ${\cal N}\sim{\cal O}(100)$ would then seem to be excluded by the observational bound, and only ${\cal N}\lesssim 10$ would be allowed. However, the actual value $\Delta N_{\rm eff}$ depends on the SM couplings in a more subtle way. To see this explicitly, consider $\cal N$} light axions $\phi_{i}$ with a coupling to QCD, so that
@@
for $c_{G}=\sqrt{\sum_{i}(c_{G}^{i})^{2}}$ and the QCD axion is identified as the linear combination $a_{\rm QCD}=(1/c_{G})\sum_{i}c_{G}^{i}\phi_{i}$. Thus, in this interaction basis, only the QCD axion interacts with the SM, while the remaining $({\cal N}-1)$ states are sterile. As a result, we would again get $\Delta N_{\rm eff}\approx 0.027$, despite having $\cal N$} axions in the spectrum. This exercise illustrates that in the presence of additional SM couplings in a scenario with $\cal N$} axions, a systematic degree of freedom counting and $\Delta N_{\rm eff}$ computation is necessary. While Ref. Gendler:2023kjt discussed axion freeze-in through their photon couplings for ${\cal N}>1$, the previous literature has primarily focused on the cosmological abundance of a single axion. For example, Refs. Cicoli:2012aq; Acharya:2015zfk; Gorbunov:2017ayg; Baer:2022fou focused on axion production via moduli decay, while the freeze-in of individual axion-like particles has been well-studied Brust:2013ova; Baumann:2016wac; Arias-Aragon:2020shv; Green:2021hjh; DEramo:2021usm; DEramo:2023asj; DEramo:2023nzt; DEramo:2024jhn.
@@
We use this EFT description to compute $\Delta N_{\rm eff}$ in the axiverse, assuming a standard radiation-dominated cosmology, with arbitrary reheating temperatures $T_{\rm RH}$ and axion decay constants $f_{a}$, and for a variety of assumptions about the origin of axion-SM couplings. We place new constraints using recent Planck Planck:2018vyg, ACT ACT:2025tim, and SPT SPT-3G:2025bzu datasets, and forecast sensitivities for Simons Observatory SimonsObservatory:2018koc; SimonsObservatory:2025wwn, CMB-S4 CMB-S4:2016ple, and CMB-HD Sehgal:2019ewc. We find that the flavor structure of the axion-fermion couplings is important in determining the observability of the cosmological axion background in the near future. Adopting a bottom-up perspective we find that existing CMB data requires that the Universe reheated to a couple orders of magnitude below $f_{a}$, or hosts very few axions. Our result emphasizes the necessity of performing top-down computations of axion-SM couplings to sharpen the predictions for $\Delta N_{\rm eff}$.
The rest of this work is structured as follows. In Sec. 2, we first review the EFT coupling of SM fields with one axion, and then generalize that discussion to $\cal N$} axions. We take into account appropriate field redefinitions to remove redundant operators. In Sec. 3, we focus on several benchmark EFTs motivated by KSVZ Kim:1979if; Shifman:1979if and DFSZ Dine:1981rt; Zhitnitsky:1980tq-type couplings. We pay attention to the flavor structure by considering anarchy Hall:1999sn, Froggatt-Neilsen textures Froggatt:1978nt, and minimal flavor violation DAmbrosio:2002vsn. In Sec. 4, we compute the freeze-in production rate of the axions. We demonstrate that for dimension-5 axion couplings, the freeze-in abundance (at temperatures much larger than the weak scale) can be computed analytically, even including the full phase space distribution of the axions. Finally, in Sec. 5, we use these rates to derive the results for $\Delta N_{\rm eff}$. We derive constraints from existing CMB data, highlight the interplay with astrophysical and flavor constraints, and identify targets for future CMB surveys. We conclude in Sec. 6.
@@ 2.1 Single Axion EFT
Before discussing an EFT containing $\cal N$} axions, it is useful to review the EFT of a single axion $\phi$ coupled to the unbroken phase of the SM Bauer:2017ris; Bauer:2020jbp; Biekotter:2023mpd; Galda:2021hbr. This EFT can most generally be written at dimension five as (in four-component Dirac fermion notation),
@@ 2.2.1 Dimension-5 EFT
where ${\cal N}_{\rm ind}$ is the number of independent axion-SM operators in Eq. (12). Note, while the mass matrix $M_{ij}$ is not diagonal in this new basis, the effect of mass mixing will not be important for the light axions considered in this work (see Sec. 4.1, and Refs. Cyncynates:2021xzw; Cyncynates:2022wlq; Cyncynates:2023esj; Chadha-Day:2023wub; Dunsky:2025sgz; Murai:2023xjn; Murai:2025wbg for scenarios where mass mixing of axions is relevant). Therefore, we do not consider the mass matrix from now on. For an explicit example of the above steps with ${\cal N}=3$ and ${\cal N}_{\rm ind}=2$, see App. A.1.
@@ 2.2.2 Dimension-6 EFT
These operators are fundamentally different from those at dimension five because they are quadratic in the axion fields. The first operator also mediates the lowest-order nonredundant coupling to the Higgs, and has previously been considered in Bauer:2022rwf. The correction analogous to (8), originating from removing the dimension-5 Higgs coupling, can be absorbed into the coefficients $c_{aH}^{ij}$. Note that the Higgs coupling is symmetric in $i\leftrightarrow j$, while the hypercharge coupling is antisymmetric. The latter class of operator, known as a charge-radius operator, has been considered before Bai:2012yq; Pierce:2014spa, but not in the context of the axiverse. Finally, replacing $B$ with $\tilde{B}$ results in an operator that vanishes upon using the Equations of Motion (EOM).
@@
Accordingly, we remain agnostic about the form of the matrix $c_{aB}^{ij}$ and just specify the vector $\lambda_{aB}^{i}$. We note that the symmetric dimension-6 operator is obtained in the IR in the KSVZ and DFSZ axion models Biekotter:2023mpd, but the skew-symmetric operator is zero if ${\cal N}=1$.
Given that the dimension-6 operators generally couple all axion states to the SM, we do not consider operators of dimension seven or higher. If dimension-6 operators are suppressed, however, dimension-7 or higher operators could be the leading production channels for some axion states, albeit more suppressed by powers of $f_{a}$. We summarize the number of independent axion states available at each dimension in Tab. 1.
@@ 2.3.1 Dirac Masses
where we work in the usual weak-interaction basis where $Y_{e}$ is diagonal. Using the total $U(1)_{L}$ symmetry we can remove the trace of ${\bf c}_{L}$, so these additional Lagrangian terms correspond to only two additional axions beyond the assumption of massless neutrinos. Finally, we note the contribution to $\Delta N_{\rm eff}$ from light right-handed neutrinos themselves is unobservably small Luo:2020fdt, unless there are additional neutrino self-interactions.
@@ 2.3.2 Majorana Masses
The simplest UV completions (e.g., see-saw Yanagida:1979as; Gell-Mann:1979vob; Glashow:1979nm; Ramond:1979py) require heavy right-handed neutrinos above the scale $\Lambda$. A careful evaluation of the associated axion production in that case is beyond the scope of this work; nevertheless, we note that our results only strictly hold for $T_{\rm RH},f_{a}<\Lambda$ if the neutrino masses are Majorana-type. Below this scale, the axion production from Eq. (29) can be ignored since the interaction is suppressed by $\Lambda f_{a}/|{\bf c}_{5}|$, as opposed to $f_{a}^{2}$, and $\Lambda/|{\bf c}_{5}|>f_{a}$ for almost all of our parameter space. Furthermore, there is a kinematic suppression due to the quintic interaction and the three-body final-state phase space. Therefore we expect that additional axion production is highly suppressed.
@@ 3.1 Hadronic Axion Couplings
In one of the most minimal scenarios, the axions do not possess tree-level couplings to the SM fermions. Such a setup is realized in the context of hadronic axion models, e.g., the KSVZ axion Kim:1979if; Shifman:1979if. A simple UV completion yielding this scenario is $\cal N$} independent copies of the KSVZ axion, such that there is a $(U(1)_{\rm PQ})^{\cal N}$ PQ symmetry which is spontaneously broken to unity at the scale ff. [^1] While this UV completion may seem less appealing when ${\cal N}\gg 1$, in theories with several compact extra dimensions, ${\cal N}\gg 1$ appears naturally thanks to the many ways of wrapping higher form fields around the compact dimensions. Below the joint PQ-breaking scale, the $\cal N$} axions acquire couplings to gauge bosons and to the Higgs through the dimension-6 operator, but not to the fermions at tree-level. The EFT can then be written as
@@
The above three axions need not be independent. For example, if the SM is embedded in a GUT with gauge group $G_{\rm GUT}$, then we get a single linear combination that has an anomalous coupling to the unified gauge group $G_{\rm GUT}$ at dimension five Srednicki:1985xd (see also Agrawal:2022lsp for a recent work):

### v2

@@ 1 Introduction
The Strong Charge-Parity (CP) problem in the Standard Model (SM) is the statement that the strong interactions appear to conserve CP symmetry while the weak interactions nearly maximally violate CP. Indeed, the Strong CP-violating parameter $\bar{\theta}$ is constrained to be $|\bar{\theta}|\lesssim 10^{-10}$ Abel and others (2020), while one would have expected $\bar{\theta}$ to be an ${\cal O}(1)$ number between $\pi$ to $\pi$, since a vanishing $\bar{\theta}$ does not lead to any enhanced symmetry of the SM. The existence of the quantum chromodynamics (QCD) axion is a hypothesized solution to this fine tuning problem. The QCD axion couples to gluons in the same way as how $\bar{\theta}$ appears in the SM Lagrangian. The presence of the QCD axion leads to a dynamical relaxation of (the effective) $\bar{\theta}$ to a vanishing value as the QCD axion gets localized to the minimum of its potential, developed during the QCD phase transition Peccei and Quinn (1977b, a); Weinberg (1978); Wilczek (1978). The QCD axion may also make up the dark matter (DM) of the Universe Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983).
@@
In minimal models, the QCD axion arises as the pseudo-Nambu-Goldstone boson of a global U(1) symmetry Kim (1979); Shifman et al. (1980); Dine et al. (1981); Zhitnitsky (1980). In such cases, the QCD axion solution to the Strong CP problem may not be robust. Planck scale-suppressed higher dimensional operators can break the global $U(1)$ symmetry and modify the QCD axion potential significantly, altering the locations of the minima of the axion potential. As a result, the QCD axion can get localized to a new minimum which does not correspond to a vanishing value of $\bar{\theta}$ and the Strong CP problem would reappear. This is known as the ‘axion quality problem’ Georgi et al. (1981); Dine and Seiberg (1986); Kamionkowski and March-Russell (1992); Holman et al. (1992); Barr and Seckel (1992); Ghigna et al. (1992), since a robust solution to the Strong CP problem demands a ‘high quality’ symmetry that is robust to various symmetry breaking effects.
@@
Models of the extra-dimensional QCD axion typically have such a high quality symmetry. A classic example of this is the 5D $U(1)$ gauge field where the axion arises from the fifth component of the gauge field. The higher dimensional gauge invariance Choi (2004); Arkani-Hamed et al. (2003), or an electric one-form symmetry of the extra-dimensional theory Reece (2025); Craig and Kongsore (2025), provides a more robust protection against symmetry breaking effects. Contributions from symmetry breaking effects lead to exponentially-small corrections, in contrast to polynomial corrections found in minimal models. This mechanism also plays out in string theory, where a large number $\cal N$} (as many as 491, found in Type IIB string theory analysis in Demirtas et al. (2020); Cicoli et al. (2012), or 181,200 in F-theory Fallon et al. (2025)) of pseudoscalars appear naturally as zero-modes of higher-dimensional gauge fields upon dimensional reduction. This leads to the notion of an ‘Axiverse’ Witten (1984); Choi and Kim (1985); Barr (1985); Svrcek and Witten (2006); Arvanitaki et al. (2010) – a universe with many axions.
@@
The QCD axion is extremely light with mass $m_{a}\sim 6\,\mu{\rm eV}\,(f_{a}/10^{12}\,{\rm GeV})$ Grilli di Cortona et al. (2016) where $f_{a}$ is the axion decay constant that controls the axion coupling with the SM, with astrophysical constraints requiring $f_{a}\gtrsim 10^{9}$ GeV Navas and others (2024). Interactions with the SM in the early universe may produce a “cosmic axion background” Turner (1987); Masso et al. (2002); Hannestad et al. (2005); Graf and Steffen (2011); Salvio et al. (2014); Ferreira and Notari (2018); Dror et al. (2021); D’Eramo et al. (2022b, 2021); Dunsky et al. (2024); Langhoff et al. (2022), which can still be around as relativistic radiation today. The axion energy density $\rho_{a}$, in comparison with the photon energy density $\rho_{\gamma}$, can be parametrized by the $\Delta N_{\rm eff}$ parameter,
@@
While the QCD axion alone contributes a small amount to $\Delta N_{\rm eff}$, one might expect the multitude of axions from the axiverse to give rise to a $\Delta N_{\rm eff}$ that is too large and ruled out by current the CMB and big-bang nucleosynthesis constraint, $\Delta N_{\rm eff}\lesssim 0.3$ Aghanim and others (2020) at 95% confidence. That is, assuming each axion attains a thermal abundance, we have $\Delta N_{\rm eff}\approx 0.027{\cal N}$. A scenario with ${\cal N}\sim{\cal O}(100)$ would then seem to be excluded by the observational bound, and only ${\cal N}\lesssim 10$ would be allowed. However, the actual value $\Delta N_{\rm eff}$ depends on the SM couplings in a more subtle way. To see this explicitly, consider $\cal N$} light axions $\phi_{i}$ with a coupling to QCD, so that
@@
for $c_{G}=\sqrt{\sum_{i}(c_{G}^{i})^{2}}$ and the QCD axion is identified as the linear combination $a_{\rm QCD}=(1/c_{G})\sum_{i}c_{G}^{i}\phi_{i}$. Thus, in this interaction basis, only the QCD axion interacts with the SM, while the remaining $({\cal N}-1)$ states are sterile. As a result, we would again get $\Delta N_{\rm eff}\approx 0.027$, despite having $\cal N$} axions in the spectrum. This exercise illustrates that in the presence of additional SM couplings in a scenario with $\cal N$} axions, a systematic degree of freedom counting and $\Delta N_{\rm eff}$ computation is necessary. While Ref. Gendler et al. (2024) discussed axion freeze-in (and decay Yin et al. (2025)) through their photon couplings for ${\cal N}>1$, the previous literature has primarily focused on the cosmological abundance of a single axion. For example, Refs. Higaki and Takahashi (2012); Cicoli et al. (2013); Higaki et al. (2013); Acharya and Pongkitivanichkul (2016); Gorbunov and Tokareva (2017); Baer et al. (2022) focused on axion production via moduli decay, while the freeze-in of individual axion-like particles both from SM couplings Brust et al. (2013); Baumann et al. (2016); Arias-Aragón et al. (2021); Green et al. (2022); D’Eramo and Yun (2022); Badziak and Laletin (2025); D’Eramo and Lenoci (2024) and from SM+BSM couplings Sakurai and Takahashi (2025); D’Eramo (2023); D’Eramo et al. (2024) has been well-studied.
@@
We use this EFT description to compute $\Delta N_{\rm eff}$ in the axiverse, assuming a standard radiation-dominated cosmology, with arbitrary reheating temperatures $T_{\rm RH}$ and axion decay constants $f_{a}$, and for a variety of assumptions about the origin of axion-SM couplings. We place new constraints using recent Planck Aghanim and others (2020), ACT Calabrese and others (2025), and SPT Camphuis and others (2025) datasets, and forecast sensitivities for Simons Observatory Ade and others (2019); Abitbol and others (2025), CMB-S4 Abazajian and others (2016), and CMB-HD Sehgal and others (2019). We find that the flavor structure of the axion-fermion couplings is important in determining the observability of the cosmological axion background in the near future. Adopting a bottom-up perspective we find that existing CMB data requires that the Universe reheated to a couple orders of magnitude below $f_{a}$, or hosts very few axions. Our result emphasizes the necessity of performing top-down computations of axion-SM couplings to sharpen the predictions for $\Delta N_{\rm eff}$.
The rest of this work is structured as follows. In Sec. 2, we first review the EFT coupling of SM fields with one axion, and then generalize that discussion to $\cal N$} axions. We take into account appropriate field redefinitions to remove redundant operators. In Sec. 3, we focus on several benchmark EFTs motivated by KSVZ Kim (1979); Shifman et al. (1980) and DFSZ Dine et al. (1981); Zhitnitsky (1980)-type couplings. We pay attention to the flavor structure by considering anarchy Hall et al. (2000), Froggatt-Nielsen textures Froggatt and Nielsen (1979), and minimal flavor violation D’Ambrosio et al. (2002). In Sec. 4, we compute the freeze-in production rate of the axions. We demonstrate that for dimension-5 axion couplings, the freeze-in abundance (at temperatures much larger than the weak scale) can be computed analytically, even including the full phase space distribution of the axions. Finally, in Sec. 5, we use these rates to derive the results for $\Delta N_{\rm eff}$. We derive constraints from existing CMB data, highlight the interplay with astrophysical and flavor constraints, and identify targets for future CMB surveys. We conclude in Sec. 6.
@@ 2.1 Single Axion EFT
Before discussing an EFT containing $\cal N$} axions, it is useful to review the EFT of a single axion $\phi$ coupled to the unbroken phase of the SM Bauer et al. (2017, 2021); Biekötter et al. (2023); Galda et al. (2021). This EFT can most generally be written at dimension five as (in four-component Dirac fermion notation),
@@ 2.2.1 Dimension-5 EFT
where ${\cal N}_{\rm ind}$ is the number of independent axion-SM operators in Eq. (12). Note, while the mass matrix $M_{ij}$ is not diagonal in this new basis, the effect of mass mixing will not be important for the light axions considered in this work (see Sec. 4.1, and Refs. Kitajima and Takahashi (2015); Daido et al. (2015, 2016); Cyncynates et al. (2022a, b); Cyncynates and Thompson (2023); Chadha-Day et al. (2024); Dunsky et al. (2025); Murai et al. (2023, 2025a, 2025b) for scenarios where mass mixing of axions is relevant). Therefore, we do not consider the mass matrix from now on. For an explicit example of the above steps with ${\cal N}=3$ and ${\cal N}_{\rm ind}=2$, see App. A.1.
@@ 2.2.2 Dimension-6 EFT
These operators are fundamentally different from those at dimension five because they are quadratic in the axion fields. The first operator also mediates the lowest-order nonredundant coupling to the Higgs, and has previously been considered in Bauer et al. (2023). The correction analogous to (8), originating from removing the dimension-5 Higgs coupling, can be absorbed into the coefficients $c_{aH}^{ij}$. Note that the Higgs coupling is symmetric in $i\leftrightarrow j$, while the hypercharge coupling is antisymmetric. The latter class of operator, known as a charge-radius operator, has been considered before Bai et al. (2013); Pierce and Zhang (2014), but not in the context of the axiverse. Finally, replacing $B$ with $\tilde{B}$ results in an operator that vanishes upon using the Equations of Motion (EOM).
@@
Accordingly, we remain agnostic about the form of the matrix $c_{aB}^{ij}$ and just specify the vector $\lambda_{aB}^{i}$. We note that the symmetric dimension-6 operator is obtained in the IR in the KSVZ and DFSZ axion models Biekötter et al. (2023), but the skew-symmetric operator is zero if ${\cal N}=1$.
Given that the dimension-6 operators generally couple all axion states to the SM, we do not consider operators of dimension seven or higher. If dimension-6 operators are suppressed, however, dimension-7 or higher operators could be the leading production channels for some axion states, albeit more suppressed by powers of $f_{a}$. We summarize the number of independent axion states available at each dimension in Tab. 1. In our results we always assume only minimal couplings to gravity, but allowing for nonminimal couplings leads to additional terms Alexander et al. (2025). However, after canonically normalizing the graviton, these terms are additionally Planck suppressed and effectively $d>6$.
@@ 2.3.1 Dirac Masses
where we work in the usual weak-interaction basis where $Y_{e}$ is diagonal. Using the total $U(1)_{L}$ symmetry we can remove the trace of ${\bf c}_{L}$, so these additional Lagrangian terms correspond to only two additional axions beyond the assumption of massless neutrinos. Finally, we note the contribution to $\Delta N_{\rm eff}$ from light right-handed neutrinos themselves is unobservably small Luo et al. (2021), unless there are additional neutrino self-interactions.
@@ 2.3.2 Majorana Masses
The simplest UV completions (e.g., see-saw Yanagida (1979); Gell-Mann et al. (1979); Glashow (1980); Ramond (1979)) require heavy right-handed neutrinos above the scale $\Lambda$. A careful evaluation of the associated axion production in that case is beyond the scope of this work; nevertheless, we note that our results only strictly hold for $T_{\rm RH},f_{a}<\Lambda$ if the neutrino masses are Majorana-type. Below this scale, the axion production from Eq. (29) can be ignored since the interaction is suppressed by $\Lambda f_{a}/|{\bf c}_{5}|$, as opposed to $f_{a}^{2}$, and $\Lambda/|{\bf c}_{5}|>f_{a}$ for almost all of our parameter space. Furthermore, there is a kinematic suppression due to the quintic interaction and the three-body final-state phase space. Therefore we expect that additional axion production is highly suppressed.
@@ 3.1 Hadronic Axion Couplings
In one of the most minimal scenarios, the axions do not possess tree-level couplings to the SM fermions. Such a setup is realized in the context of hadronic axion models, e.g., the KSVZ axion Kim (1979); Shifman et al. (1980). A simple UV completion yielding this scenario is $\cal N$} independent copies of the KSVZ axion, such that there is a $(U(1)_{\rm PQ})^{\cal N}$ PQ symmetry which is spontaneously broken to unity at the scale ff. [^1] While this UV completion may seem less appealing when ${\cal N}\gg 1$, in theories with several compact extra dimensions, ${\cal N}\gg 1$ appears naturally thanks to the many ways of wrapping higher form fields around the compact dimensions. Below the joint PQ-breaking scale, the $\cal N$} axions acquire couplings to gauge bosons and to the Higgs through the dimension-6 operator, but not to the fermions at tree-level. The EFT can then be written as
@@
The above three axions need not be independent. For example, if the SM is embedded in a GUT with gauge group $G_{\rm GUT}$, then we get a single linear combination that has an anomalous coupling to the unified gauge group $G_{\rm GUT}$ at dimension five Srednicki (1985) (see also Agrawal et al. (2022) for a recent work):